DIRC Operations Weekly Report
Period: Wednesday 2007/2/21 - Wednesday 2007/2/28
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070301.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070221
PEP peak lumi: ~8.2 x 1033
(previous week: ~6.5 x 1033)
BaBar integrated lumi: ~3.2
(previous week: ~1.7 fb-1)
Run 6 PEP delivery:~3.3
(previous week: ~1.8 fb-1)
Run 6 total delivery: 9.9 fb-1
Run 6 total BaBar recorded: 9.6 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
Foreseen work opportunities (PEP down periods):
BaBar activities/issues this week:
logbook for details
- Replace two DFBs in sector 1 with innefficient or flaky channels (require bridge).
- HV group #11 in sector 0 is noisy: needs to be masked.
- Unmask all masked channels to find which of these are actually noisy. After December DFB work/movement masked channels may no longer be the noisy ones.
Bradley, Jose and Alejandro rotating backup & on-call commissioning duties